### v1

# Egocentric Action-aware Inertial Localization in Point Clouds

Abstract
This paper presents a novel inertial localization framework named Egocentric Action-aware Inertial Localization (EAIL), which leverages egocentric action cues from head-mounted IMU signals to localize the target individual within a 3D point cloud. Human inertial localization is challenging due to IMU sensor noise that causes trajectory drift over time. The diversity of human actions further complicates IMU signal processing by introducing various motion patterns. Nevertheless, we observe that some actions observed through the head-mounted IMU correlate with spatial environmental structures (e.g., bending down to look inside an oven, washing dishes next to a sink), thereby serving as spatial anchors to compensate for the localization drift. The proposed EAIL framework learns such correlations via hierarchical multi-modal alignment. By assuming that the 3D point cloud of the environment is available, it contrastively learns modality encoders that align short-term egocentric action cues in IMU signals with local environmental features in the point cloud. These encoders are then used in reasoning the IMU data and the point cloud over time and space to perform inertial localization. Interestingly, these encoders can further be utilized to recognize the corresponding sequence of actions as a by-product. Extensive experiments demonstrate the effectiveness of the proposed framework over state-of-the-art inertial localization and inertial action recognition baselines. Project page: https://github.com/mf-zhang/Ego-Inertial-Localization.
1 Introduction
Figure 1: Egocentric Action-aware Inertial Localization (EAIL). Our framework leverages egocentric action cues obtained from the head-mounted IMU to perform inertial localization in the environmental 3D point cloud. The corresponding sequence of actions can also be recognized as a by-product.
Inertial localization has emerged as a pivotal technology for human tracking across a wide range of applications, including personal navigation [22] and augmented reality [52]. Inertial measurement units (IMUs) are commonly embedded in wearable devices such as smartphones, smart glasses [12], and headsets [46]. When affixed to human bodies, IMUs can capture acceleration and angular velocity to record 3D human movements. Compared to vision-based localization methods [40, 28], inertial localization enables user tracking in an energy-efficient and privacy-preserving manner [30].
Despite the advantages, human inertial localization remains highly challenging primarily due to two reasons. The first one is the trajectory drift caused by IMU noise. IMU sensor noise causes small uncertainties in measurements that accumulate over time, which can ultimately lead to significant trajectory drift [59]. Conventional step detection methods [3, 65] struggle to generalize to the noise of irregular movements, while recent data-driven approaches [22, 67, 55, 38], which predict velocity from IMU signals, can suffer from accumulated estimation errors.
The other challenge lies in the complexity of human actions. Human body-mounted IMUs capture signals not only from displacement-related movements like walking and stopping, but also from action-induced motions that do not involve actual positional changes, such as head-swinging during cooking and cleaning. These additional motion signals can complicate IMU signal processing and make inertial localization further difficult. Existing datasets [63, 22, 5] and approaches [24, 63] focus mostly on human walking scenarios, limiting the ability of state-of-the-art inertial localization methods to handle real-world human action variability.
Nevertheless, we argue that human actions can rather act as a salient locational cue to mitigate the trajectory drift challenge if properly taken into account. For example, washing dishes often takes place near a sink, while frying often occurs near a stove. Our key insight is that extracting such actions from IMU data can help to deduce the spatial regions where they are likely to occur.
In this work, we present a novel framework named Egocentric Action-aware Inertial Localization (EAIL; see also Fig. 1). While assuming that the 3D point cloud of the environment is readily available using off-the-shelf 3D scanners, this framework leverages egocentric action cues extracted using head-mounted IMUs to localize the target individual in the environmental point cloud.
The key technical contribution is a hierarchical modality alignment technique that effectively extracts the correlation between the action cues and the point cloud. Specifically, we first learn the encoders for short-term IMU signals and local point clouds to be aligned with the corresponding egocentric point-of-view images and textual descriptions of observed actions in a contrastive fashion. The learned IMU and point-cloud encoders are then invoked in spatio-temporal reasoning modules to jointly predict possible actions occurring at each moment and the locations in the global point cloud where the predicted actions can be observed. While our primary goal is the localization in the latter part, the predicted locations can also improve inertial action recognition as a beneficial by-product.
Extensive evaluations on the EgoExo4D dataset [18] validate that our framework achieves state-of-the-art performance in both inertial localization and inertial action recognition compared to [24, 67, 42, 70]. In summary, our main contributions are as follows:
  • We introduce EAIL, a novel inertial localization framework that leverages egocentric action cues from head-mounted IMU signals to localize target individuals within a 3D point cloud.
  • We develop a hierarchical modality alignment technique that learns the correlation between the inertial egocentric action cues and the environmental point cloud to perform inertial localization as our primary goal, and action recognition as a by-product.
  • Extensive evaluations demonstrate that our framework achieves state-of-the-art results in both inertial localization and inertial action recognition in diverse settings.
2 Related Work
2.1 Human Inertial Localization
Human inertial localization has been an active research area. Leveraging both visual and inertial data, accurate localization can be achieved by perceiving and modeling the surrounding environment [43, 32, 50]. Visual-inertial fusion can also simultaneously perform localization and body poses estimation [36, 64]. However, visual data availability is not guaranteed at all times, especially in human-centric applications where privacy concerns are essential.
Traditional inertial localization methods like Pedestrian Dead Reckoning (PDR) [54, 3, 65] estimate trajectories by detecting steps and estimating step length and heading. While effective in structured environments, these approaches struggle with noise from irregular human motion. Data-driven inertial navigation methods [4, 67, 22] mitigate sensor noise by learning to estimate velocity from IMU signals. However, when they cumulate the estimated velocity to predict human locations, the estimation errors can lead to significant trajectory drift.
To mitigate the drift problem, external signals such as GPS [29], Wi-Fi [23], Bluetooth [33], user history [66], and activities [13, 14] have been used to provide anchor points for human localization. However, these solutions introduce dependencies on specific infrastructures or hand-defined landmarks. More recently, NILoc [24] takes a different approach by directly predicting user locations without explicit velocity integration, but it requires scene-specific training and lacks adaptability across diverse environments.
Finally, beyond these methodological challenges, existing datasets [22, 5] and approaches [24, 63] are primarily centered around displacement-centric activities such as walking and stopping. This focus overlooks a broader spectrum of complex human actions, thus limiting the models’ abilities to manage real-world human motion variability. In contrast, our framework addresses the trajectory drift problem by leveraging egocentric action cues from IMU signals, to localize the user within a point cloud.
2.2 Egocentric Multimodal Alignment
Recent advancements in head-mounted devices [12, 52] have generated significant interest in understanding human activities from an egocentric perspective. Several large-scale egocentric datasets [17, 18, 8, 25] have been introduced, offering rich multimodal data including video, gaze, audio, language, and IMU signals. These datasets provide a foundation for developing multimodal human understanding methods [16, 42, 37, 35].
Aligning egocentric visual data with textual action descriptions enhances comprehension and broadens application possibilities. Large-scale video-language pretraining, specifically tailored to egocentric data, has been pioneered by works like EgoVLP [34, 48] and HierVL [2]. In addition to the human language narrations, LaViLa [71] refines these narrations with Large Language Models (LLMs). Recap [27] and Vinci [26] align fused languages for long videos. EMBED [11] and EgoInstructor [61] use rules or retrieval models to add additional training data. Subsequent works further incorporate additional egocentric-specific cues such as hand [68] and audio [6]. These approaches have demonstrated that egocentric-specific pretraining yields transferable representations for various downstream tasks, such as question answering [15], robot manipulation [47], and narration-based segmentation [53].
Aligning vision and language embeddings with 3D environments [62, 39, 69] makes open-vocabulary affordance detection possible. Affordance denotes potential actions in the environment that an agent can perform to an object or in an area. [9] provides accurate interaction annotations for real-world 3D indoor scenes. Works like [56, 19, 41] utilize free-form queries describing geometry and affordances, leveraging the pretrained CLIP [51] model, to retrieve 3D instances. These studies show that a strong correlation between actions and environments can be learned through multimodal alignment.
Inspired by these advances, our framework introduces learning the multimodal alignment between the human action cues and the environmental structures, to serve as natural anchors in the scene to predict human trajectories.
3 Problem Setting
Figure 2: Short-Term Action-Location Alignment. In this first stage, our objective is to train a point cloud encoder and an IMU encoder using contrastive learning. A positive sample set consists of simultaneous multimodal data samples at the human’s location. Negative samples are generated by randomly sampling data from other times and locations.
We address the task of inertial localization using head-mounted devices equipped with IMUs. We assume that the 3D point cloud of the environment is available using consumer-grade 3D scanner apps [31, 1]. Our primary goal is to predict the sequence of the device user’s locations in the point cloud.
[…]
Note that this problem setup is different from existing inertial navigation (e.g., [22]) and inertial localization [24]. Inertial navigation methods just predict velocity and thus require a known initial position to perform localization, making them prone to cumulative drift. Existing inertial localization [24] operates in a scene-specific manner, training and evaluating models separately for each environment while disregarding explicit environmental structure. In contrast, our approach incorporates the 3D point cloud $\boldsymbol{\mathcal{P}}$, enabling localization without requiring environment-specific training.
4 The EAIL Framework
The proposed EAIL framework leverages action cues from head-mounted IMU signals to localize target individuals within the 3D point cloud. It proceeds in two stages:
  • Stage 1: Short-term action-location alignment (Fig. 2). We first train an IMU encoder for processing short-term IMU signals (e.g., spanning 1 second) as well as a point-cloud encoder for local point clouds (e.g., 1 m2). Here we propose a novel modality alignment technique that aligns the features of those IMU and point cloud inputs with the corresponding egocentric point-of-view images and text annotations of the actions taken by the target individual, which makes the trained encoders ‘action-aware’. Note that these egocentric images and text annotations are necessary only in this training stage.
  • Stage 2: Sequential motion localization (Fig. 3). We then learn spatiotemporal reasoning modules that invoke the IMU/point-cloud encoders to predict the sequence of target individual’s locations (and the sequence of their actions as a bonus; see Sec. 4.2).
4.1 Stage 1: Short-term Action-Location Alignment
In the first stage, we aim to train a short-term IMU signal encoder and a local point cloud encoder. In the training, we focus on capturing action semantics to establish a meaningful correspondence between the two modalities. These correlations, such as the association between dish-washing motion and the structure of a sink, can serve as natural anchors to enhance downstream localization accuracy.
To achieve this goal, we employ multimodal contrastive learning. Egocentric datasets [18, 17, 8] usually include vision and text data alongside IMU data. We conduct contrastive alignment across four modalities: images, textual action descriptions, IMU signals, and point clouds, as shown in Fig. 2. In this way, we can utilize the robust capabilities of pre-trained and pre-aligned vision-language models [44, 51, 7]. The action semantics embedded in vision and text modalities aid to guide the training of the IMU and the point cloud encoders. Note that the image and the language modalities are not required during testing.
Figure 3: Sequential Motion Localization. In this second stage, we generate a sequence of the user’s locations and actions over $T$ seconds using a series of IMU signals ($\{\textbf{M}_{t}\}_{t=1}^{T}$) alongside the point cloud of the entire scene. This point cloud is divided into $S$ local segments ($\{\textbf{P}_{s}\}_{s=1}^{S}$). The IMU encoder and the point cloud encoder are frozen networks pre-trained in Stage 1 for efficient spatial-temporal reasoning in large-scale point clouds.
Specifically, for each 1-second time segment, we extract a synchronized set of inputs: an ego-view image frame $\textbf{I}_{t}$, an action caption $\textbf{L}_{t}$, an IMU segment $\textbf{M}_{t}$, and a local $1m^{2}$ region of point cloud ${\textbf{P}}_{t}$. We utilize off-the-shelf pre-trained vision-language encoders $E_{I}$ and $E_{L}$ to encode the image and the text data, while we train the IMU and the point cloud encoders, $E_{M}$ and $E_{P}$. Then, we apply a pairwise contrastive loss $L_{c}$ across the generated features $\textbf{F}^{I}$, $\textbf{F}^{L}$, $\textbf{F}^{M}$, and $\textbf{F}^{P}$ derived from the four modalities:
$\begin{split}L_{stage1}=\alpha L_{c}(\textbf{F}^{I},\textbf{F}^{M})+\beta L_{c}(\textbf{F}^{I},\textbf{F}^{P})\\ +\theta L_{c}(\textbf{F}^{L},\textbf{F}^{M})+\delta L_{c}(\textbf{F}^{L},\textbf{F}^{P})+\gamma L_{c}(\textbf{F}^{M},\textbf{F}^{P}).\end{split}$ (1)
Through this process, our IMU and point cloud encoders are trained to generate features in a shared embedding space where multimodal features are connected through human action semantics. This enables the retrieval of local point cloud segments with short-term IMU signals. It can provide a preliminary estimate of the user’s location with IMU signals. However, given that similar motion patterns might occur in multiple locations, we introduce the second stage of our framework to address these ambiguities.
4.2 Stage 2: Sequential Motion Localization
In Stage 2, we build on the pre-trained short-term IMU encoder and the local point cloud encoder, and invoke them for processing sequential motion captured over $T$ seconds of IMU signals $\boldsymbol{\mathcal{M}}$ and the global point cloud $\boldsymbol{\mathcal{P}}$ of the entire scene to predict the user’s trajectory $\boldsymbol{\mathcal{Z}}$.
4.2.1 Spatiotemporal reasoning for trajectory prediction
Predicting user trajectories within large-scale point clouds requires fine-grained temporal and spatial reasoning. To achieve this, as shown in Fig. 3, we decompose the $T$-second IMU signals $\boldsymbol{\mathcal{M}}$ into short sub-sequences, $\{\textbf{M}_{t}\}_{t=1}^{T}$. Concurrently, the global point cloud is uniformly partitioned into $S$ local segments, denoted as $\{\textbf{P}_{s}\}_{s=1}^{S}$. Then, we apply the frozen $E_{M}$ to extract IMU features $\{\textbf{F}^{M}_{t}\}_{t=1}^{T}$. In parallel, the frozen $E_{P}$ processes local point cloud segments to produce point cloud features $\{\textbf{F}^{P}_{s}\}_{s=1}^{S}$.
Given that IMU features $\textbf{F}^{M}$ and point cloud features $\textbf{F}^{P}$ are aligned due to the Stage 1 contrastive learning process, we can generate a sequence of correspondence heatmaps $\{\textbf{H}_{t}\}_{t=1}^{T}$ by calculating the similarity between these two feature sets. Regions with a high score on these heatmaps indicate a high likelihood that human motion occurs in these spatial regions. For instance, “washing” would align closely with a “sink” region, while “walking” covers “open areas”.
Building on this, we design a Temporal Reasoning Module and a Spatial Reasoning Module. The temporal reasoning module takes the correspondence heatmaps $\{\textbf{H}_{t}\}_{t=1}^{T}$ and IMU features $\{\textbf{F}^{M}_{t}\}_{t=1}^{T}$ as inputs and employs a 3D convolutional network to reason across the temporal dimension, generating refined features $\textbf{F}^{R}$. Next, $\textbf{F}^{R}$ is inputted, along with the point cloud features $\{\textbf{F}^{P}_{s}\}_{s=1}^{S}$, into the spatial reasoning module. This module uses another dilated 3D convolutional network to reason in the 2D spatial dimension. The final output is the predicted user trajectory $\boldsymbol{\mathcal{Z}}$.
For the trajectory $\boldsymbol{\mathcal{Z}}$, we frame the prediction as an $S$-class classification problem. The goal is to find a sequence of point cloud segments that are nearest to the actual user trajectory. We use cross-entropy loss for training:
$L_{traj}=-\sum_{t=1}^{T}\sum_{s=1}^{S}\textbf{y}_{t,s}\log(\hat{\textbf{y}}_{t,s}),$ (2)
where $\textbf{y}_{t,s}$ is the true label indicating the user’s nearest point cloud segment and $\hat{\textbf{y}}_{t,s}$ is the predicted probability for time step $t$ and segment $P_{s}$.
4.2.2 Location-aware action recognition
In the Location-aware Action Recognition Module, we leverage the predicted user locations to enhance inertial action recognition. Specifically, for each time step $t$, we take all the predicted location probabilities $\hat{\textbf{y}}_{t,s}$ in a scene, collectively represented by heatmaps $\{\textbf{H}^{\prime}_{t}\}_{t=1}^{T}$, and use them as a spatial attention over the point cloud features $\{\textbf{F}^{P}_{s}\}_{s=1}^{S}$. We then blend these spatial features with IMU features $\{\textbf{F}^{M}_{t}\}_{t=1}^{T}$ through addition. Finally, a multi-layer perceptron maps the fused representation to action likelihood. The training is supervised by a cross-entropy loss:
$L_{action}=-\sum_{t=1}^{T}\sum_{c\in\mathcal{C}}\boldsymbol{y}_{t,c}\log(\hat{\boldsymbol{y}}_{t,c}),$ (3)
where $\boldsymbol{y}_{t,c}$ represents the true label for the action class $c$ at time $t$, and $\hat{\boldsymbol{y}}_{t,c}$ is the predicted probability.
In summary, our whole model in Stage 2 is supervised by $L_{stage2}=L_{traj}+L_{action}$.
5 Experiments
5.1 Experimental Setup
Dataset
We use the EgoExo4D dataset [18] that provides synchronized egocentric video and IMU signals recorded with Aria glasses [12]. Prior to data collection, environment point clouds were built, and action captions were annotated afterward. For this paper, we used the cooking activities subset, which includes 173 participants across 60 kitchens, totaling 564.13 hours of recordings. The activity area for these cooking activities averaged around 2.8 meters per side, with the largest spanning 6.15 meters. Each local 1 $m^{2}$ point cloud is sub-sampled to contain 8192 points, and the IMU signals are recorded at an 800 Hz sample rate. The dataset is labeled with 35 distinct action classes for classification tasks. We divided the data into training, test-seen, and test-unseen sets. The test-unseen set consists exclusively of environments and participants not present in the training set.
Implementation Details
In Stage 1, we employed an enhanced version of the CLIP model [44] as the vision-language encoder. Specifically, we used ViT-Base [10] as the image encoder and the Base Text Transformer [60] from CLIP for text encoding. For the point cloud encoder, we adopted PointNet++ [49], and the IMU encoder was built upon ResNet18-1D [20]. The IMU signals are pre-processed following [22] and downsampled to 400 Hz. In Stage 2, our model operated with a sequence length $T=10$, uniformly partitioning the point cloud into $S=20\times 20=400$ local segments. We trained Stage 1 for 250 epochs and Stage 2 for 100 epochs, using a batch size of 64, a learning rate of $10^{-3}$, and the AdamW optimizer. The loss parameters were set as follows: $\alpha=0.1;\ \beta,\theta,\delta,\gamma=1$.
Evaluation Metrics
For the localization task, we report the success rate (%) at error distance thresholds of 0.2 m, 0.4 m, and 0.6 m following NILoc [24]. Additionally, we include the Relative Score (RS) metric, which measures the proportion of heatmap positions with lower confidence than the ground truth location. This metric indicates the relative ranking of the prediction at the target point. For action classification, we evaluate performance using top-1 and top-5 accuracy metrics. These scores are assessed under two setups: “seen rooms” where the localization is performed in the environments present in the training dataset and “unseen rooms” where environments are otherwise new.
5.2 Inertial Localization Results
Baselines
RoNIN [22] learns to predict velocity from IMU signals. We select ResNet-50 as the most effective backbone. IMUNet [67] introduces a new architecture and implements it with various backbones [58, 57]. NILoc [24] directly predicts user location with IMU signals. We choose to use ResNet-18 to encode IMU singals to maintain consistency with our framework. We train all the aforementioned methods on the EgoExo4D dataset.
Velocity Cumulation versus Direct Location Prediction
Previous approaches can be divided into two categories. The first category starts with a given initial localization, estimates velocity using IMU signals, and accumulates the velocity to obtain the target’s location, including [22, 58, 57, 67]. The second category directly predicts the target’s location, including [24] and ours. As shown in Tab. 1, direct location prediction methods outperform velocity accumulation methods. This is because each recording in the dataset is relatively long, with an average duration of 525 seconds and a maximum of 2,526 seconds. Velocity accumulation methods are significantly affected by error accumulation over such long durations.
Table 1: Inertial Localization Results. We evaluate the accuracy using two metrics: the localization success rate (%) at various error distance thresholds and the Relative Score (RS) metric for localization likelihood prediction (methods that do not generate likelihood predictions are not evaluated with RS). For both metrics, higher values indicate better performance.
{tblr}
colspec=X[1,l] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c], cell11 = r=2l, cell13 = c=4c, cell18 = c=4c, hline1,9 = -0.08em, hline2 = 3-6, hline2 = 8-11, hline3 = -, width=1.1 Method & Seen Rooms Unseen Rooms
0.2m 0.4m 0.6m RS 0.2m 0.4m 0.6m RS
MnasNet[58] 2.95 7.57 12.75 / 2.16 5.45 9.86 /
EfficientNet[57] 3.1 7.77 12.71 / 2.54 6.68 10.84 /
IMUNet[67] 3.74 10.15 17.23 / 3.4 9.17 14.63 /
RoNIN[22] 4.86 12.77 20.65 / 3.96 9.52 15.65 /
NILoc[24]$+$ 17.03 41.31 74.15 88.17 13.32 37.85 69.21 84.08
Ours 43.86 70.15 89.6 96.01 26.86 65.97 90.79 89.55
Figure 4: Inertial Localization Error Over Time Elapsed. While velocity accumulation-based methods experience significant trajectory drift, our approach remains accurate over time.
Trajectory Drift
Fig. 4 illustrates the difference between the two categories of methods with the change in average position error over elapsed time. In this figure, velocity accumulation methods require a given initial position, resulting in zero error at time zero. However, due to error accumulation over time, their error increases rapidly, surpassing that of our method at around 30 seconds. In contrast, direct localization prediction methods do not require an initial position and maintain a stable error over time.
Direct Localization in Point Clouds
NILoc [24] and our method both directly predict a user’s location from IMU signals. However, NILoc relies solely on inertial data, so it can only fit a single scene at a time. In their original work, they trained three separate models, each for a different scene. For a fair comparison, we developed NILoc+, which uses the same multi-scene training data as our Stage 2. Nevertheless, its lack of spatial awareness still leads to reduced accuracy, whereas our approach leverages point cloud structures to deliver robust inertial localization across diverse environments.
5.3 Inertial Action Recognition Results
Baselines
DeepConvLSTM [45] uses convolutional networks and LSTMs to classify actions from IMU signals. EVIMAE [70] leverages the MAE framework [21] to learn representations from both video and multiple IMU signals for action classification. For comparison with our approach, we train a modified version of EVIMAE using data from a single IMU device, employing ViT-Base as the backbone. IMU2CLIP [42] uses a strategy similar to our Stage 1, employing a pretrained CLIP model [51, 44] to guide IMU feature extraction and fine-tuning with an MLP head for action recognition. For a fair comparison, we use the same backbone for each modality and the same training data as our method when training IMU2CLIP.
Table 2: Inertial Action Recognition Results. We evaluate performance using top1 and top5 accuracy metrics. Higher values indicate better performance.
{tblr}
colspec=X[1,l] X[0.1,c] X[1,c] X[1,c] X[0.1,c] X[1,c] X[1,c], cell11 = r=2c, cell13 = c=2c, cell16 = c=2c, hline1,7 = -0.08em, hline2 = 3-4, hline2 = 6-7, hline3 = -, width=1.25 Method & Seen Rooms Unseen Rooms
top1 top5 top1 top5
DeepConvLSTM[45] 15.20 43.27 12.47 36.86
EVIMAE[70]-- 18.34 46.12 9.30 31.10
IMU2CLIP[42] 18.96 50.43 12.27 37.04
Ours 21.48 53.62 15.03 43.34
Quantitative Evaluation
As shown in Tab. 2, DeepConvLSTM performs relatively poorly due to its architecture’s difficulty in capturing complex action patterns effectively. Although EVIMAE gains from the advanced reasoning capabilities inherent in transformer networks, it still faces challenges when generalizing to unseen rooms. Much like our method, IMU2CLIP leverages the learning of action semantics from vision-language modalities; however, it doesn’t quite match the effectiveness of our approach. This is largely because our method incorporates spatial reasoning, utilizing environmental information to enhance action recognition performance.
Table 3: Ablation Studies. We report the inertial localization accuracy and the inertial action recognition ($\mathcal{A}$) accuracy simultaneously.
{tblr}
colspec=X[2.25,l] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c], cell11 = r=2l, cell13 = c=6c, cell110 = c=6c, cell31 = c=15l,fg=myblue!50, cell61 = c=15l,fg=myblue!50, cell91 = c=15l,fg=myblue!50, hline1 = -0.08em, hline2 = 3-8, hline2 = 10-15, hline3 = -0.08em, hline6 = -0.08em, hline9 = -0.08em, hline13 = -0.08em, width=1.2 Method & Seen Rooms Unseen Rooms
0.2m 0.4m 0.6m RS $\mathcal{A}$-top1 $\mathcal{A}$-top5 0.2m 0.4m 0.6m RS $\mathcal{A}$-top1 $\mathcal{A}$-top5
Modalities Engagement (visual and textual action description benefits localization)
w/o vision language 39.75 66.30 87.56 95.70 13.21 51.02 20.41 61.95 89.83 88.13 9.53 29.09
w/ vision language 43.86 70.15 89.6 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
Action Classification Task (explicit action supervision benefits localization)
w/o action loss 41.92 68.54 87.75 95.51 / / 25.37 63.79 89.28 89.02 / /
w/ action loss 43.86 70.15 89.6 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
Spatial and Temporal Reasoning
w/o spatial 38.68 66.13 87.78 95.05 21.20 52.96 25.03 59.39 83.54 88.56 14.40 43.77
w/o temporal 41.44 68.48 87.96 95.78 21.22 52.48 26.41 64.15 89.13 89.41 14.61 43.06
w/ both 43.86 70.15 89.6 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
Figure 5: Visualization of Heatmaps in Each Stage.
5.4 Ablation Studies
Modalities Engagement for Action-aware Alignment
In Stage 1 of our framework, we focus on effectively training the IMU and the point cloud encoders to yield features that are aligned through action semantics. To accomplish this, we leverage the power of robust pre-trained and pre-aligned vision-language models, such as [51, 44]. These models excel at capturing intricate associations between visual content and corresponding textual descriptions. Our experiments, as detailed in Tab. 3, affirm the efficacy of incorporating vision and language modalities. By doing so, our model is able to learn and align features across multiple modalities, significantly enhancing both localization and action recognition accuracy. Importantly, it’s worth noting that while the integration of vision and language modalities provides substantial benefits during the training phase, they are not required during the inference phase. Furthermore, even in scenarios where action caption annotations are unavailable in the training set, our method does not fail, ensuring reasonable accuracy without relying on complete annotation sets. This demonstrates the broad applicability and flexibility of our approach in real-world environments.
Action Classification Supervision in Stage 2
We incorporate action classification supervision in Stage 2. The results in Tab. 3 show that including the action loss leads to more accurate location predictions. We believe that explicitly learning action categories helps the model align human motion more effectively with the surrounding environment.
Table 4: Location-Aware Action Recognition Ablation Study. “PC” denotes point cloud features, and “LA” represents location attention.
{tblr}
colspec=X[1,l] X[0.1,c] X[1,c] X[1,c] X[0.1,c] X[1,c] X[1,c], cell11 = r=2c, cell13 = c=2c, cell16 = c=2c, hline1,6 = -0.08em, hline2 = 3-4, hline2 = 6-7, hline3 = -, width=1.25 Method & Seen Rooms Unseen Rooms
top1 top5 top1 top5
w/o PC 18.81 50.23 12.37 38.21
w/o LA 17.88 46.22 10.43 35.63
w/ LA PC 21.48 53.62 15.03 43.34
Spatial and Temporal Reasoning in Stage 2
In the second stage of our framework, we leverage a temporal reasoning module for comprehending a sequence of IMU signals, and a spatial reasoning module for understanding the global environmental point cloud. This dual-layered reasoning framework is essential, as evidenced in Tab. 3. It shows that when these two modules work in tandem, they align sequential motion data with the environmental settings. This holistic reasoning capability ensures the model generates a more precise and coherent trajectory prediction of the user’s movements within their environment.
Spatial Attention Benefits Inertial Action Recognition
After obtaining the inertial localization results, we find that incorporating surrounding environmental structure cues is beneficial for inertial action recognition. As shown in Tab. 4, using only IMU signals, we achieve results comparable to IMU2CLIP [42]. Naively incorporating global point cloud features with IMU features leads to a performance drop. In contrast, integrating point cloud features with predicted location attention with IMU features provides a clear performance improvement.
More Ablation Results in Supplementary Material
Further ablation results can be found in Tab. 5 in our Supplementary Material. These include analyses on different vision-language encoders in Stage 1, the preliminary location retrieval accuracy in Stage 1, different architecture designs in Stage 2, and different choices of sequence temporal length in Stage 2.
5.5 Qualitative evaluations
Visualization of Heatmaps in Each Stage
Our framework is capable of producing highly interpretable intermediate results in the form of heatmaps. The heatmap from Stage 1 reflects the direct similarity strength between the features generated by the IMU encoder and the point cloud encoder. The heatmap from Stage 2 represents the localization likelihood at each location, considering both the motion sequence and the global scene structure. As shown in Fig. 5, in Kitchen 1, two peak points appear in the Stage 1 heatmap, indicating that the motion in the IMU signals could plausibly occur at either of these locations. However, after the spatiotemporal reasoning in Stage 2, our framework successfully identifies a single, distinct peak.
Comparison with Previous Methods
In Fig. 6, we present a qualitative comparison with RoNIN[22] and NILoc[24]+. The recorded sequence features a person taking 12 minutes to make a cup of coffee. From the visualization results, we observe that RoNIN suffers from cumulative errors, causing its predicted location to drift outside the point cloud boundary after just 2 minutes. NILoc+, on the other hand, lacks spatial awareness and may produce predictions in physically implausible locations, such as areas already occupied by objects. In contrast, our method leverages action cues such as the upward head motion when fetching items from a cabinet, to establish action-environment correlations, enabling accurate localization.
Figure 6: Qualitative Comparison with Previous Methods.
6 Limitations and Future Directions
While our method can robustly exploit head-mounted IMU signals for human localization within pre-built point clouds, it does hinge on several factors that present avenues for future research. First, the requirement of a complete and up-to-date 3D scan limits applicability in environments subject to frequent layout changes, motivating efforts on incremental or online map updates. Second, our reliance on action-environment correlations makes localization challenging when users remain idle (e.g., standing still for extended periods). Integrating standard inertial navigation solutions and Kalman filters could help bridge gaps during low-motion segments. Finally, our experiments are based on IMU data from head-mounted devices, and substantially different sensor placements (e.g., ankle or wrist) may necessitate model adaptations for robust performance.
7 Conclusion
We introduce a novel framework EAIL for inertial localization that uniquely leverages egocentric action cues from head-mounted IMU signals to enhance localization within 3D point clouds. By aligning short-term IMU signals with local point cloud structures, and further integrating temporal and spatial reasoning, our approach effectively addresses common challenges like trajectory drift and complex human actions. Additionally, our method demonstrates a robust capacity for both trajectory prediction and action recognition across diverse settings, as validated by rigorous evaluations on the EgoExo4D dataset. This work underscores the potential for action-environment correlations to serve as powerful anchors in practical inertial localization tasks, opening avenues for future explorations into more advanced sensing and positioning technologies.
Supplementary Material
Table 5: Supplementary Ablation Studies.
{tblr}
colspec=X[2.25,l] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c], cell11 = r=2l, cell13 = c=6c, cell110 = c=6c, cell31 = c=15l,fg=myblue!50, cell71 = c=15l,fg=myblue!50, cell111 = c=15l,fg=myblue!50, cell151 = c=15l,fg=myblue!50, hline1 = -0.08em, hline2 = 3-8, hline2 = 10-15, hline3 = -0.08em, hline7 = -0.08em, hline11 = -0.08em, hline15 = -0.08em, hline18 = -0.08em, width=1.1 Method & Seen Rooms Unseen Rooms
0.2m 0.4m 0.6m RS $\mathcal{A}$-top1 $\mathcal{A}$-top5 0.2m 0.4m 0.6m RS $\mathcal{A}$-top1 $\mathcal{A}$-top5
Different Vision-Language Encoder in Stage 1
CLIP 42.72 69.33 89.63 95.94 20.81 52.05 21.60 53.45 83.41 86.44 12.68 40.38
InternVL 42.41 69.21 90.95 96.15 17.15 50.80 26.11 60.46 87.70 90.18 9.84 39.44
SLIP 43.86 70.15 89.6 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
Temporal Length of Input IMU Sequence in Stage 2
5s 45.27 71.86 90.84 96.18 21.23 53.57 25.69 64.07 87.11 89.28 14.53 43.62
20s 43.10 69.61 89.11 95.84 20.82 52.53 24.77 61.09 85.99 89.15 15.41 44.77
10s 43.86 70.15 89.6 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
Residual Connection in Stage 2 Architecture Design
w/o IMU residual 36.29 63.56 86.84 95.05 21.45 52.38 26.07 60.41 86.03 89.10 14.90 43.15
w/o PC residual 39.14 64.71 86.16 95.30 20.24 52.08 23.51 59.66 86.02 87.78 14.60 43.77
w/ both 43.86 70.15 89.6 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
Preliminary Location Retrieval Accuracy in Stage 1
Stage 1 Retrieval 18.29 41.23 66.70 85.47 / / 14.25 38.51 64.31 78.31 / /
Full 2 Stages 43.86 70.15 89.6 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
8 Supplementary Ablation Studies
We provide additional ablation results shown in Tab. 5.
Different Vision-Language Encoder in Stage 1
We compare CLIP-Large[51], InternVL[7], and SLIP-Base[44] as possible backbones for the vision-language encoders. While InternVL scores the highest on some metrics, SLIP generally attains stronger localization accuracy. Hence, in our default configuration, we adopt SLIP as the vision-language encoder.
Temporal Length of Input IMU Sequence in Stage 2
We experiment with sequences of 5s, 10s, and 20s of IMU data. Shorter 5s windows excel on the seen rooms but underperform on the unseen set. Longer 20s windows better preserve temporal context and yield a slight gain in unseen action top‐1 performance, yet they degrade localization in both settings. By contrast, 10s serves as a balanced choice, delivering robust localization and action recognition across seen/unseen scenarios.
Residual Connection in Stage 2 Architecture Design
Removing residual connections, either on the IMU feature path or on the point cloud feature path, consistently degrades results. In contrast, preserving both forms of residual connections significantly improves localization and action accuracy. This highlights the importance of allowing the network to fuse new spatiotemporal cues while maintaining a direct pathway for unimodal features.
Preliminary Location Retrieval Accuracy in Stage 1
Lastly, we assess how well the short-term feature alignment from Stage 1 alone can recover the user’s location. Doing so attains considerably lower accuracy compared to our full Stage 2 inference. This underscores the utility of leveraging both temporal context and global spatial reasoning to refine the initial retrieval from Stage 1.
9 Supplementary Video
We create a supplementary video that further illustrates the performance of our framework. The heatmaps featured in the video are interpreted similarly to those shown in Fig. 5, providing visual insight into the localization process. The video demonstrates the robustness of our approach, highlighting its capability for accurate long-term human tracking across extended periods.

### v2

@@
-Egocentric Action-aware Inertial Localization in Point Clouds
+Egocentric Action-aware Inertial Localization in Point Clouds with Vision-Language Guidance
 Abstract
@@
 The diversity of human actions further complicates IMU signal processing by introducing various motion patterns.
-Nevertheless, we observe that some actions observed through the head-mounted IMU correlate with spatial environmental structures (e.g., bending down to look inside an oven, washing dishes next to a sink), thereby serving as spatial anchors to compensate for the localization drift.
-The proposed EAIL framework learns such correlations via hierarchical multi-modal alignment.
+Nevertheless, we observe that some actions captured by the head-mounted IMU correlate with spatial environmental structures (e.g., bending down to look inside an oven, washing dishes next to a sink), thereby serving as spatial anchors to compensate for the localization drift.
+The proposed EAIL framework learns such correlations via hierarchical multi-modal alignment with vision-language guidance.
 By assuming that the 3D point cloud of the environment is available, it contrastively learns modality encoders that align short-term egocentric action cues in IMU signals with local environmental features in the point cloud.
-These encoders are then used in reasoning the IMU data and the point cloud over time and space to perform inertial localization.
+The learning process is enhanced using concurrently collected vision and language signals to improve multimodal alignment.
+The learned encoders are then used in reasoning the IMU data and the point cloud over time and space to perform inertial localization.
 Interestingly, these encoders can further be utilized to recognize the corresponding sequence of actions as a by-product.
@@
 The corresponding sequence of actions can also be recognized as a by-product.
-Inertial localization has emerged as a pivotal technology for human tracking across a wide range of applications, including personal navigation [22] and augmented reality [52].
+Inertial localization has emerged as a pivotal technology for human tracking across a wide range of various applications, including personal navigation [22] and augmented reality [51].
 Inertial measurement units (IMUs) are commonly embedded in wearable devices such as smartphones, smart glasses [12], and headsets [45].
 When affixed to human bodies, IMUs can capture acceleration and angular velocity to record 3D human movements.
-Compared to vision-based localization methods [40, 28], inertial localization enables user tracking in an energy-efficient and privacy-preserving manner [30].
+Compared to vision-based localization methods [39, 28], inertial localization enables user tracking in an energy-efficient and privacy-preserving manner.
 Despite the advantages, human inertial localization remains highly challenging primarily due to two reasons.
@@
 While assuming that the 3D point cloud of the environment is readily available using off-the-shelf 3D scanners, this framework leverages egocentric action cues extracted using head-mounted IMUs to localize the target individual in the environmental point cloud.
-The key technical contribution is a hierarchical modality alignment technique that effectively extracts the correlation between the action cues and the point cloud.
+The key technical contribution is a hierarchical modality alignment technique that effectively extracts the correlation between the action cues and the point cloud with vision-language guidance.
 Specifically, we first learn the encoders for short-term IMU signals and local point clouds to be aligned with the corresponding egocentric point-of-view images and textual descriptions of observed actions in a contrastive fashion.
+By incorporating guidance from pretrained vision-language models, we achieve improved multimodal feature alignment.
 The learned IMU and point-cloud encoders are then invoked in spatio-temporal reasoning modules to jointly predict possible actions occurring at each moment and the locations in the global point cloud where the predicted actions can be observed.
@@
 For both metrics, higher values indicate better performance.
-{tblr}
-colspec=X[1,l] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c], cell11 = r=2l, cell13 = c=4c, cell18 = c=4c, hline1,9 = -0.08em, hline2 = 3-6, hline2 = 8-11, hline3 = -, width=1.1 Method & Seen Rooms Unseen Rooms
+Method $SeenRooms$ Unseen Rooms
 $0.2m$ 0.4m 0.6m RS 0.2m 0.4m 0.6m RS
-MnasNet[58] 2.95 7.57 12.75 / 2.16 5.45 9.86 /
-EfficientNet[57] 3.1 7.77 12.71 / 2.54 6.68 10.84 /
-IMUNet[67] 3.74 10.15 17.23 / 3.4 9.17 14.63 /
-RoNIN[22] 4.86 12.77 20.65 / 3.96 9.52 15.65 /
-NILoc[24]$+$ 17.03 41.31 74.15 88.17 13.32 37.85 69.21 84.08
-Ours 43.86 70.15 89.6 96.01 26.86 65.97 90.79 89.55
+MnasNet[57] 2.952.952.95 7.57 12.75 / 2.16 5.45 9.86 /
+EfficientNet[56] 3.103.103.10 7.77 12.71 / 2.54 6.68 10.84 /
+IMUNet[66] 3.743.743.74 10.15 17.23 / 3.40 9.17 14.63 /
+RoNIN[22] 4.864.864.86 12.77 20.65 / 3.96 9.52 15.65 /
+NILoc[24]$+$ 17.0317.0317.03 41.31 74.15 88.17 13.32 37.85 69.21 84.08
+Ours 43.86 70.15 89.60 96.01 26.86 65.97 90.79 89.55
 Figure 4: Inertial Localization Error Over Time Elapsed. While velocity accumulation-based methods experience significant trajectory drift, our approach remains accurate over time.
@@
 Table 2: Inertial Action Recognition Results. We evaluate performance using top1 and top5 accuracy metrics. Higher values indicate better performance.
-{tblr}
-colspec=X[1,l] X[0.1,c] X[1,c] X[1,c] X[0.1,c] X[1,c] X[1,c], cell11 = r=2c, cell13 = c=2c, cell16 = c=2c, hline1,7 = -0.08em, hline2 = 3-4, hline2 = 6-7, hline3 = -, width=1.25 Method & Seen Rooms Unseen Rooms
+Method $SeenRooms$ Unseen Rooms
 $top1$ top5 top1 top5
-DeepConvLSTM[45] 15.20 43.27 12.47 36.86
-EVIMAE[70]-- 18.34 46.12 9.30 31.10
-IMU2CLIP[42] 18.96 50.43 12.27 37.04
+DeepConvLSTM[44] 15.2015.2015.20 43.27 12.47 36.86
+EVIMAE[69]-- 18.3418.3418.34 46.12 9.30 31.10
+IMU2CLIP[41] 18.9618.9618.96 50.43 12.27 37.04
 Ours 21.48 53.62 15.03 43.34
@@
 Table 3: Ablation Studies. We report the inertial localization accuracy and the inertial action recognition ($\mathcal{A}$) accuracy simultaneously.
-{tblr}
-colspec=X[2.25,l] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c], cell11 = r=2l, cell13 = c=6c, cell110 = c=6c, cell31 = c=15l,fg=myblue!50, cell61 = c=15l,fg=myblue!50, cell91 = c=15l,fg=myblue!50, hline1 = -0.08em, hline2 = 3-8, hline2 = 10-15, hline3 = -0.08em, hline6 = -0.08em, hline9 = -0.08em, hline13 = -0.08em, width=1.2 Method & Seen Rooms Unseen Rooms
+Method $SeenRooms$ Unseen Rooms
 $0.2m$ 0.4m 0.6m RS $\mathcal{A}$-top1 $\mathcal{A}$-top5 0.2m 0.4m 0.6m RS $\mathcal{A}$-top1 $\mathcal{A}$-top5
 Modalities Engagement (visual and textual action description benefits localization)
-w/o vision language 39.75 66.30 87.56 95.70 13.21 51.02 20.41 61.95 89.83 88.13 9.53 29.09
-w/ vision language 43.86 70.15 89.6 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
+w/o vision language 39.7539.7539.75 66.30 87.56 95.70 13.21 51.02 20.41 61.95 89.83 88.13 9.53 29.09
+w/ vision language 43.86 70.15 89.60 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
 Action Classification Task (explicit action supervision benefits localization)
-w/o action loss 41.92 68.54 87.75 95.51 / / 25.37 63.79 89.28 89.02 / /
-w/ action loss 43.86 70.15 89.6 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
+w/o action loss 41.9241.9241.92 68.54 87.75 95.51 / / 25.37 63.79 89.28 89.02 / /
+w/ action loss 43.86 70.15 89.60 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
 Spatial and Temporal Reasoning
-w/o spatial 38.68 66.13 87.78 95.05 21.20 52.96 25.03 59.39 83.54 88.56 14.40 43.77
-w/o temporal 41.44 68.48 87.96 95.78 21.22 52.48 26.41 64.15 89.13 89.41 14.61 43.06
-w/ both 43.86 70.15 89.6 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
+w/o spatial 38.6838.6838.68 66.13 87.78 95.05 21.20 52.96 25.03 59.39 83.54 88.56 14.40 43.77
+w/o temporal 41.4441.4441.44 68.48 87.96 95.78 21.22 52.48 26.41 64.15 89.13 89.41 14.61 43.06
+w/ both 43.86 70.15 89.60 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
 Figure 5: Visualization of Heatmaps in Each Stage.
@@
 Table 4: Location-Aware Action Recognition Ablation Study. “PC” denotes point cloud features, and “LA” represents location attention.
-{tblr}
-colspec=X[1,l] X[0.1,c] X[1,c] X[1,c] X[0.1,c] X[1,c] X[1,c], cell11 = r=2c, cell13 = c=2c, cell16 = c=2c, hline1,6 = -0.08em, hline2 = 3-4, hline2 = 6-7, hline3 = -, width=1.25 Method & Seen Rooms Unseen Rooms
+Method $SeenRooms$ Unseen Rooms
 $top1$ top5 top1 top5
-w/o PC 18.81 50.23 12.37 38.21
-w/o LA 17.88 46.22 10.43 35.63
+w/o PC 18.8118.8118.81 50.23 12.37 38.21
+w/o LA 17.8817.8817.88 46.22 10.43 35.63
 w/ LA PC 21.48 53.62 15.03 43.34
@@
 7 Conclusion
-We introduce a novel framework EAIL for inertial localization that uniquely leverages egocentric action cues from head-mounted IMU signals to enhance localization within 3D point clouds.
-By aligning short-term IMU signals with local point cloud structures, and further integrating temporal and spatial reasoning, our approach effectively addresses common challenges like trajectory drift and complex human actions.
-Additionally, our method demonstrates a robust capacity for both trajectory prediction and action recognition across diverse settings, as validated by rigorous evaluations on the EgoExo4D dataset.
-This work underscores the potential for action-environment correlations to serve as powerful anchors in practical inertial localization tasks, opening avenues for future explorations into more advanced sensing and positioning technologies.
+We propose EAIL, a novel framework for inertial localization that uses egocentric action cues from head-mounted IMU signals to improve positioning in 3D point clouds.
+By aligning short-term IMU data with local structures and incorporating temporal-spatial reasoning, EAIL addresses challenges like trajectory drift and complex actions.
+Evaluated on the EgoExo4D dataset, it shows strong performance in trajectory prediction and action recognition.
+This highlights the potential of action-environment correlations as anchors for inertial localization, paving the way for advanced sensing and positioning methods.
+8 Acknowledgement
+This work was supported by JSPS KAKENHI Grant Number JP24K02956, JST ASPIRE Grant Number JPMJAP2303, JST SPRING Grant Number JPMJSP2108.
 Supplementary Material
 Table 5: Supplementary Ablation Studies.
-{tblr}
-colspec=X[2.25,l] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[0.1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c] X[1,c], cell11 = r=2l, cell13 = c=6c, cell110 = c=6c, cell31 = c=15l,fg=myblue!50, cell71 = c=15l,fg=myblue!50, cell111 = c=15l,fg=myblue!50, cell151 = c=15l,fg=myblue!50, hline1 = -0.08em, hline2 = 3-8, hline2 = 10-15, hline3 = -0.08em, hline7 = -0.08em, hline11 = -0.08em, hline15 = -0.08em, hline18 = -0.08em, width=1.1 Method & Seen Rooms Unseen Rooms
+Method $SeenRooms$ Unseen Rooms
 $0.2m$ 0.4m 0.6m RS $\mathcal{A}$-top1 $\mathcal{A}$-top5 0.2m 0.4m 0.6m RS $\mathcal{A}$-top1 $\mathcal{A}$-top5
 Different Vision-Language Encoder in Stage 1
-CLIP 42.72 69.33 89.63 95.94 20.81 52.05 21.60 53.45 83.41 86.44 12.68 40.38
-InternVL 42.41 69.21 90.95 96.15 17.15 50.80 26.11 60.46 87.70 90.18 9.84 39.44
+CLIP 42.7242.7242.72 69.33 89.63 95.94 20.81 52.05 21.60 53.45 83.41 86.44 12.68 40.38
+InternVL 42.4142.4142.41 69.21 90.95 96.15 17.15 50.80 26.11 60.46 87.70 90.18 9.84 39.44
 SLIP 43.86 70.15 89.6 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
@@
 5s 45.27 71.86 90.84 96.18 21.23 53.57 25.69 64.07 87.11 89.28 14.53 43.62
-20s 43.10 69.61 89.11 95.84 20.82 52.53 24.77 61.09 85.99 89.15 15.41 44.77
-10s 43.86 70.15 89.6 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
+20s 43.1043.1043.10 69.61 89.11 95.84 20.82 52.53 24.77 61.09 85.99 89.15 15.41 44.77
+10s 43.8643.8643.86 70.15 89.6 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
 Residual Connection in Stage 2 Architecture Design
-w/o IMU residual 36.29 63.56 86.84 95.05 21.45 52.38 26.07 60.41 86.03 89.10 14.90 43.15
-w/o PC residual 39.14 64.71 86.16 95.30 20.24 52.08 23.51 59.66 86.02 87.78 14.60 43.77
+w/o IMU residual 36.2936.2936.29 63.56 86.84 95.05 21.45 52.38 26.07 60.41 86.03 89.10 14.90 43.15
+w/o PC residual 39.1439.1439.14 64.71 86.16 95.30 20.24 52.08 23.51 59.66 86.02 87.78 14.60 43.77
 w/ both 43.86 70.15 89.6 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
 Preliminary Location Retrieval Accuracy in Stage 1
-Stage 1 Retrieval 18.29 41.23 66.70 85.47 / / 14.25 38.51 64.31 78.31 / /
+Stage 1 Retrieval 18.2918.2918.29 41.23 66.70 85.47 / / 14.25 38.51 64.31 78.31 / /
 Full 2 Stages 43.86 70.15 89.6 96.01 21.48 53.62 26.86 65.97 90.79 89.55 15.03 43.34
 9 Supplementary Ablation Studies
-We provide additional ablation results shown in Tab. 5.
 Different Vision-Language Encoder in Stage 1